# Membrane phononic integrated circuits

### v1

@@ V.1 Justification of assumptions
The harmonic balance analysis assumed that the quality factor is large and that the Duffing nonlinearity is relatively weaker than the linear terms. The quality factor assumption can be justified knowing that for materials such as silicon nitride that have low intrinsic dissipation, low clamping losses, and a high stress such that dissipation dilution further increases the quality factor [14, 15], $Q\gg 1$ is easily achievable for most resonator geometries. Even at room temperature and atmospheric pressure, small resonators of comparable size to the atmospheric mean free path length can achieve significant quality factors up to $Q\simeq 10^{2}-10^{3}$ [79, 80, 81].
@@
$\frac{\text{nonlinear forces}}{\text{linear forces}}$ $\leq\frac{Y}{2\sigma}\left[\max\left(\left(\frac{\partial u}{\partial x}\right)^{2}\right)+\max\left(\left(\frac{\partial u}{\partial y}\right)^{2}\right)\right]$
$=\frac{Y}{2\sigma}\times 2\left(\frac{\pi Z_{0}}{W}\right)^{2}$
@@ V.2 Bistability
We saw in Fig. 12 that at large enough values of $\alpha_{\mathrm{eff}}$ the original Lorentzian response curve tilts over and becomes multivalued, accepting three distinct amplitude solutions. The middle amplitude is not experimentally realisable because it corresponds to a saddle point of the dynamical system [76]. However, the other two amplitudes are stable solutions. When driven at a frequency at which bistability occurs, the resonator can respond at either of those two different amplitudes of motion. Which amplitude it responds at depends on the system history; in other words it exhibits hysteresis. [76, 24, 78].
To understand how this works in practice, imagine holding the drive amplitude constant and let $\alpha_{\mathrm{eff}}>0$, so we follow the orange curve in Fig. 12. As $\Omega_{d}$ is swept from 0 to $\infty$, the response amplitude will slowly increase until it achieves its peak value, then sudden ‘jump down’ to the lower branch. If we perform the opposite sweep from $\infty$ to 0, the drive will smoothly move along the lower branch until the point marked by the red cross, then ‘jump up’ to the higher branch. On the green curve where $\alpha_{\mathrm{eff}}<0$ the same jumps will occur but during the opposite sweeps.
@@
Figure 13: Bistability of a Duffing oscillator. Black line: analytic solution from solving Eq. (44), showing stable solutions (solid) and also unstable solutions (dashed). Markers: solutions from numerically solving Eq. (42). Red arrows: increasing drive amplitude with ‘jump up’ transition. Blue arrows: decreasing amplitude sweep with ‘jump down’ transition. The drive force $F$ and displacement AA are in arbitrary units.
@@ V.3 Critical amplitude
The critical amplitude unsurprisingly decreases when $\sqrt{\alpha_{\mathrm{eff}}/k_{\mathrm{eff}}}$, the ratio of nonlinear to linear force per unit displacement, increases. It also decreases when $Q$ is larger, because narrower resonance peaks require smaller frequency shifts to reach bistability (see Fig. 12(b)).
@@ V.4 Geometric Duffing nonlinearity
As said at the beginning of Section V, there are two key sources of Duffing nonlinearity for suspended membrane phononics: elongation stress in the material during vibration, and electrostatic forces. In this section explain the origin of elongation stresses and derive an expression for the corresponding Duffing coefficient.
At small scales gravity is a far weaker force than the internal stresses of a membrane. For example, consider a silicon nitride membrane with density $\rho=3200\,\mathrm{kg\,m^{-3}}$ and tensile stress $\sigma=1\,\mathrm{GPa}$. In one dimension, when hanging under its own weight, it assumes a catenary (hyperbolic cosine) curve with minimum radius of curvature $\sigma/g\rho\approx 32\,\mathrm{km}$ [90]. This is vastly greater than the membrane’s horizontal scale, so we can safely say the membrane is flat when at rest [91, 92]. Therefore, because the edges of the membrane are fixed, any curvature of the membrane (such as from vibration) must require some elongating strain in the material. This is the root cause of what is called the geometric Duffing nonlinearity [24].
@@
Writing $u(x)=Z_{0}\psi(x)$ as the product of the maximum displacement and a modeshape $\psi(x)\in[-1,1]$ shows us the strain energy is proportional to the fourth power of $Z_{0}$:
@@
As an example, consider Fig. 15, where we use finite element software (COMSOL Multiphysics [42]) to model a pair of silicon nitride resonators, one a concave two-dimensional beam, and the other a doubly-tethered trampoline. Both resonators effectively share the same eigenfrequency around $3.8\,\mathrm{MHz}$. However, their effective Duffing coefficients are almost two orders of magnitude ($77\times$) different. This is because, as seen in Fig. 15(b), the design of the trampoline concentrates the flexural strain to the tethers. The tethers are clamped to the (horizontal) substrate, so motion there necessarily requires curvature of the material, leading to geometric nonlinearity. The concavity of the beam does the opposite effect, creating a more even distribution of curvature throughout the resonator. This is to the technique of soft clamping [16], which minimises material curvature in order to reduce quality factors.
@@ V.5 Electrostatic actuation
In addition to geometric elongation, a second common source of Duffing nonlinearity in membrane phononics is electrostatic forces. Here we explain the principle of electrostatic actuation, derive the corresponding Duffing coefficient, and explain other nuances such as linear resonance frequency shift and the existence of quadratic nonlinearity.
@@
Membranes can be electrostatically actuated by depositing an electrode on the membrane surface, then creating a potential difference between that electrode and another fixed electrode. For example, one can fabricate gold wiring on top of the membrane via a metal liftoff process, then apply a voltage to the wires while grounding the substrate. Alternatively, the substrate could be replaced with a suspended electrode above the membrane [12]. These two setups are illustrated in Fig. 16(a) and (b) respectively.
@@
Consider the more common situation where the substrate is used as the fixed electrode. The other electrode is patterned onto the membrane and has area $A$ and voltage $V$. Because the etch depth $d$ ($\approx 500\,\mathrm{nm}$) is typically much less than the electrode size ($\approx 10\mathrm{\upmu m}$), and considering the extremely large ($\approx 10^{4})$ megahertz-range dielectric permittivity of doped silicon [20], the electric field is well-localised to the vertical column beneath the electrode as in an ideal parallel plate capacitor. Indeed finite element simulations show that for typical values the effective capacitance is only half a percentage point off the value for a parallel plate capacitor [20].
@@
We can understand the dynamical effect of this force by Taylor expanding it about $z/d$:
@@
This force corresponds to a static displacement of the membrane towards the substrate, which we can find by considering the Duffing equation (Eq. (42)) for a constant displacement $z(t)=Z_{0}$:
@@
Here $k_{\mathrm{eff}}$ is the effective spring constant determined by the material and geometry using Eq. (26). Assuming small deflections such that $k_{\mathrm{eff}}\gg\alpha_{\mathrm{eff}}Z_{0}^{2}$, the new equilibrium displacement will be shifted by $\Delta Z_{0}\simeq F_{\mathrm{cap},0}/k_{\mathrm{eff}}$. For a linear resonator this displacement is all that occurs and the dynamical behaviour is unchanged. However, if we also factor in the Duffing nonlinearity from Section V.4 we expect the longitudinal strain from the displacement to modify the effective spring constant. The value of the change can be obtained by taking the derivative of the left hand side of Eq. (59), recognising that the spring constant is the force proportional to the displacement. The change in spring constant is:
@@
In this case the term is negative (softening) so it acts to decrease the resonance frequency. A DC voltage can be used for this reason to decrease the eigenfrequency of a nanomechanical oscillator [98, 99]. Comparing the spring constant with and without the electrostatic force, we can see that the resonance frequency shifts from $\Omega_{0}=\sqrt{k_{\mathrm{eff}}/m_{\mathrm{eff}}}$ to:
@@
Whether the change in frequency is positive or negative depends on the applied voltage and the etch depth. $\Delta k_{\mathrm{Duff}}$ has the same sign as $\alpha_{\mathrm{eff}}$, which for most materials including silicon nitride is positive, and scales with $V^{4}/d^{4}$. $\Delta k_{\mathrm{cap}}$ is negative and scales with $V^{2}/d^{3}$. The different scaling laws imply the frequency shift can be positive or negative depending on $V$ and $d$, as illustrated in Fig. 17. It can be seen that the stiffening correction dominates for large voltages and the softening term dominates for small voltages (especially for smaller etch depths). Setting $\Delta k_{\mathrm{Duff}}=\Delta k_{\mathrm{cap}}$, we find that the transition occurs at $V=2k_{\mathrm{eff}}\sqrt{d/3\varepsilon_{0}\varepsilon_{r}A\alpha_{\mathrm{eff}}}$.
@@
Figure 18(b) shows that this correction not change the qualitative behaviour but does increase the magnitude of the electrostatic nonlinearity.
@@ V.6 DC bias for oscillatory motion
The first term provides a constant force that changes the effective spring constant as described in Eq. (62). The middle term provides the actuation at the desired input frequency. In experiments it is often the case that $V_{\mathrm{DC}}>V_{\mathrm{AC}}$, in which case the actuation force has been amplified beyond what could be achieved with an alternating potential alone. The last term can often be ignored, both because $V_{\mathrm{DC}}>V_{\mathrm{AC}}$ and because it is not a force at the frequency of interest.
@@ V.7.1 Pull-in
Membranes with larger actuation electrodes and smaller capacitive gaps are more susceptible to pull-in. For example, halving the electrode length or radius doubles the safe voltage range. Examples of pull-in voltages for a typical silicon nitride membrane are given in Fig. 19.
Figure 19: Equilibrium displacement calculated from Eq. (66) for different electrode areas. Blue line: electrode area of $A=10\times 10\,\mathrm{\mu m}^{2}$. Orange: $A=20\times 20\,\mathrm{\mu m}^{2}$. Green: $A=30\times 30\,\mathrm{\mu m}^{2}$. Circles mark the pull-in instability locations. The etch depth is $d=500\,\mathrm{nm}$. The effective spring constant was calculated analytically from Eq. (26) assuming a $30\,\mathrm{\mu}m$ silicon nitride square membrane vibrating in the fundamental mode with $\sigma=1\,\mathrm{GPa}$ and $\rho=3200\,\mathrm{kg\cdot m^{-3}}$.
@@ VI Evanescent fields
So far in this paper we have separately introduced membrane waveguides and resonators, along with the continuum and lumped-element models. We now shift focus towards how these circuit elements can be coupled. Coupling is an essential ingredient in phononic circuits; we will show that one of the outstanding aspects of suspended membranes is that the coupling can be engineered very flexibly and precisely. We will build towards this conclusion in the rest of the paper and finish by outlining several examples. Firstly, here we introduce evanescent acoustic fields, which underpin the mechanism by which acoustic phonons are coupled.
@@
Evanescent waves are also very useful for membrane phononic circuits, where they allow single mode waveguides and highly efficient coupling of heterogeneous membrane geometries [3, 22, 53]. A simple use case for these functions is the coupling of a resonator with input and output single mode waveguides, allowing mechanical logic to be performed without electromechanical actuation [20]. In this tutorial we will go further and detail several more highly non-trivial phononic devices based that can be achieved using evanescent couplers (see Section XI).
To see how evanescent waves arise in membrane phononics, consider a waveguide of width $W$ defined by having a density $\rho_{1}$ greater than the density $\rho_{2}$ of the surrounding material, as illustrated in Fig. 20. This is not the same mechanism of operation as suspended membrane waveguides, which are defined instead by the impedance mismatch between the membrane and substrate, but it has the same principle and makes for a simpler example.
@@
$\psi(x)=\begin{cases}\cos(k_{x}x)&|x|<\frac{W}{2}\\ e^{-\kappa x}&\text{else}.\end{cases}$ (70)
@@
$\frac{\rho_{2}\Omega^{2}}{\sigma}$ $=k_{y}^{2}-\alpha^{2}.$ (73)
Combining Eq.s (71), (72), and (73) eliminates $k_{y}$ and $\kappa$, yielding a transcendental equation for $k_{x}$:
@@
Figure 21: Equivalence between confinement of optical waves with dielectric index contrast and confinement of acoustic waves with density contrast. (a) In a dielectric slab or optical fiber, the guided mode is confined into the core, which has high refractive index than cladding material. The mode evanescently extends into the cladding, but no energy is transferred outside so there is total internal reflection (dashed line). (b) A mathematically equivalent example is the phononic waveguide studied in this section, where density differences create a guided mode.
Equation (70) is mathematically equivalent to the case of the transverse electric mode of light in a dielectric waveguide, if we make the substitution $\rho\mapsto\varepsilon$ [56]; that is, if we replace density differences with electric permittivity differences. This equivalence is illustrated in Fig. 21. Because refractive index scales with $\varepsilon^{2}$, the density contrast of $\rho_{1}/\rho_{2}=4$ illustrated in Fig. 20 is equivalent to a refractive index contrast of $n_{2}/n_{1}=2$, such as is found for silicon nitride photonic waveguides surrounded by air [118].

### v2

@@ V.1 Justification of assumptions
The harmonic balance analysis assumed that the quality factor is large and that the Duffing nonlinearity is relatively weaker than the linear terms. The quality factor assumption can be justified knowing that for materials such as silicon nitride (which have low intrinsic dissipation, low clamping losses, and a high stress such that dissipation dilution further increases the quality factor [14, 15]) $Q\gg 1$ is easily achievable for most resonator geometries. Even at room temperature and atmospheric pressure, small resonators of comparable size to the atmospheric mean free path length can achieve significant quality factors up to $Q\simeq 10^{2}-10^{3}$ [87, 88, 89].
@@
$\frac{\text{nonlinear forces}}{\text{linear forces}}$ $\leq\frac{Y}{2\sigma}\left[\max\left(\left(\frac{\partial u}{\partial x}\right)^{2}+\left(\frac{\partial u}{\partial y}\right)^{2}\right)\right]$
$\leq\frac{Y}{2\sigma}\times 2\left(\frac{\pi Z_{0}}{W}\right)^{2}$
@@ V.2 Bistability
We saw in Fig. 12 that at large enough values of $\alpha_{\mathrm{eff}}$ the original Lorentzian response curve tilts over and becomes multivalued, accepting three distinct amplitude solutions. The middle amplitude is not experimentally realisable because it corresponds to a saddle point of the dynamical system [84]. However, the other two amplitudes are stable solutions. When driven at a frequency at which bistability occurs, the resonator can respond at either of those two different amplitudes of motion. Which amplitude it responds at depends on the system history; in other words it exhibits hysteresis [84, 24, 86].
To understand how this works in practice, imagine holding the drive amplitude constant and let $\alpha_{\mathrm{eff}}>0$, so we follow the orange curve in Fig. 12. As $\Omega_{d}$ is swept from 0 to $\infty$, the response amplitude will slowly increase until it achieves its peak value, then sudden ‘jump down’ to the lower branch. If we perform the opposite sweep from $\infty$ to 0, the drive will smoothly move along the lower branch until the point marked by the red cross, then ‘jump up’ to the higher branch. On the green curve where $\alpha_{\mathrm{eff}}<0$, the same jumps will occur but during the opposite sweeps.
@@
Figure 13: Bistability of a Duffing oscillator. Black line: analytic solution from solving Eq. (45), showing stable solutions (solid) and also unstable solutions (dashed). Markers: solutions obtained by numerically solving Eq. (43). Red arrows: increasing drive amplitude with ‘jump up’ transition. Blue arrows: decreasing amplitude sweep with ‘jump down’ transition. The drive force $F$ and displacement AA are in arbitrary units.
@@ V.3 Critical amplitude
When $\sqrt{\alpha_{\mathrm{eff}}/k_{\mathrm{eff}}}$, the ratio of nonlinear to linear spring constants, increases, the critical amplitude decreases. It also decreases when $Q$ is larger, because narrower resonance peaks require smaller frequency shifts to reach bistability (see Fig. 12(b)).
@@ V.4 Geometric Duffing nonlinearity
As stated at the beginning of Section V, there are two key sources of Duffing nonlinearity for suspended membrane devices: elongation stress in the material during vibration, and electrostatic forces. In this section we explain the origin of elongation stresses and derive an expression for the corresponding Duffing coefficient.
At small scales gravity is a far weaker force than the internal stresses of a membrane. For example, consider a silicon nitride membrane with density $\rho=3200\,\mathrm{kg\cdot m^{-3}}$ and tensile stress $\sigma=1\,\mathrm{GPa}$. In one dimension, when hanging under its own weight, it assumes a catenary (hyperbolic cosine) curve with minimum radius of curvature $\sigma/(g\rho)\approx 32\,\mathrm{km}$ [98]. This is vastly greater than the membrane’s horizontal scale, so we can safely say the membrane is flat when at rest [99, 100]. Therefore, because the edges of the membrane are fixed, any curvature of the membrane (such as from vibration) must require some elongating strain in the material. This is the root cause of what is called the geometric Duffing nonlinearity [24].
@@
Writing $u(x)=Z_{0}\psi(x)$ as the product of the maximum displacement and a modeshape $\psi(x)$ where $\max(|\psi(x)|)=1$ shows us the strain energy is proportional to the fourth power of $Z_{0}$:
@@
As an example, consider Fig. 15, where we use finite element software (COMSOL Multiphysics [42]) to model a pair of silicon nitride resonators (a), one a concave two-dimensional beam, and the other a doubly-tethered trampoline (b), with both resonators drawn to the same scale. Both resonators effectively share the same eigenfrequency around $3.8\,\mathrm{MHz}$. However, their effective Duffing coefficients are almost two orders of magnitude ($77\times$) different. This is because, as seen in Fig. 15(b), the design of the trampoline concentrates the flexural strain to the tethers. The tethers are clamped to the substrate, so motion there necessarily requires a large localised deformation of the material, leading to geometric nonlinearity. The concave beam (a) produces the opposite effect, creating a more even distribution of strain throughout the resonator. This is reminiscent to the technique of soft clamping [16], which minimises material curvature in order to improve quality factors.
@@ V.5 Electrostatic actuation
In addition to geometric elongation, a second common source of Duffing nonlinearity in membrane phononics is electrostatic forces. Here we explain the principle of electrostatic actuation, derive the corresponding Duffing coefficient, and explain other nuances such as linear resonance frequency shifts and the existence of quadratic nonlinearity.
@@
Membranes can be electrostatically actuated by depositing an electrode on the membrane surface, then creating a potential difference between that electrode and another fixed electrode. For example, one can fabricate gold wiring on top of the membrane via a metal liftoff process, then apply a voltage to the wires while grounding the electrically conductive substrate [20]. Alternatively, the substrate could be replaced with a suspended electrode above the membrane [12]. These two setups are illustrated in Fig. 16(a) and (b) respectively.
@@
Consider the more common situation where the substrate is used as the fixed electrode [9, 20]. The second electrode is patterned onto the membrane and has area $A$ and voltage $V$. Because the etch depth $d$ ($\approx 500\,\mathrm{nm}$) is typically much less than the electrode size ($\approx 10\mathrm{\upmu m}$), and considering the extremely large ($\approx 10^{4})$ megahertz-range dielectric permittivity of doped silicon [20], the electric field is well-localised to the vertical column of depth $d$ beneath the electrode, as in an ideal parallel plate capacitor. Indeed finite element simulations show that for typical values the effective capacitance is only half a percentage point off the value for a parallel plate capacitor [20].
@@
We can understand the dynamical effect of this force by performing a Taylor expansion it about $z/d$:
@@
This force produces a to a static displacement of the membrane towards the substrate, which we can find by considering the Duffing equation (Eq. (43)) for a constant displacement $z(t)=Z_{0}$:
@@
Here $k_{\mathrm{eff}}$ is the effective spring constant determined by the material and geometry using Eq. (26). Assuming small deflections such that $k_{\mathrm{eff}}\gg\alpha_{\mathrm{eff}}Z_{0}^{2}$, the new equilibrium displacement will be shifted by $\Delta Z_{0}\simeq F_{\mathrm{cap},0}/k_{\mathrm{eff}}$. For a linear resonator this displacement is all that occurs and the dynamical behaviour is unchanged. However, if we also factor in the Duffing nonlinearity from Section V.4 we expect the longitudinal strain from the displacement to modify the effective spring constant. By linearising Eq. (60) around the displacement of $\Delta Z_{0}\simeq F_{\mathrm{cap},0}/k_{\mathrm{eff}}$, the change in spring constant can be obtained as:
@@
In this case the term is negative (softening) so it acts to decrease the resonance frequency. A DC voltage can be used for this reason to decrease the eigenfrequency of a nanomechanical oscillator—this is the well known process of electrostatic softening [106, 107]. Comparing the spring constant with and without the electrostatic force, we can see that the resonance frequency shifts from $\Omega_{0}=\sqrt{k_{\mathrm{eff}}/m_{\mathrm{eff}}}$ to:
@@
Whether the change in frequency is positive or negative depends on the applied voltage and the etch depth. $\Delta k_{\mathrm{Duff}}$ is positive and scales with $V^{4}/d^{4}$. $\Delta k_{\mathrm{cap}}$ is negative and scales with $V^{2}/d^{3}$. The different scaling laws imply the frequency shift can be positive or negative depending on $V$ and $d$, as illustrated in Fig. 17. It can be seen that the stiffening correction dominates for large voltages and the softening term dominates for small voltages (especially for smaller etch depths). Setting $\Delta k_{\mathrm{Duff}}=\Delta k_{\mathrm{cap}}$, we find that the transition occurs at $V=2k_{\mathrm{eff}}\sqrt{d/3\varepsilon_{0}\varepsilon_{r}A\alpha_{\mathrm{eff}}}$.
@@
Figure 18(a) shows that this correction does not change the qualitative behaviour but does increase the magnitude of the electrostatic nonlinearity.
@@ V.6 DC bias for oscillatory motion
The first term provides a constant force that changes the membrane equilibrium position and the effective spring constant as described in Eq. (63). The middle term provides the actuation at the desired input frequency. In experiments it is often the case that $V_{\mathrm{DC}}\gg V_{\mathrm{AC}}$, in which case the actuation force has been amplified beyond what could be achieved with an alternating potential alone. The last term can often be ignored, both because $V_{\mathrm{DC}}\gg V_{\mathrm{AC}}$ and because it is not a force at the frequency of interest.
@@ V.7.1 Pull-in
Membranes with larger actuation electrodes and smaller capacitive gaps are more susceptible to pull-in. Examples of pull-in voltages for a typical silicon nitride membrane are given in Fig. 19.
Figure 19: Equilibrium displacement magnitude $|z|$ calculated from Eq. (67) for different electrode areas. Blue line: electrode area of $A=10\times 10\,\mathrm{\mu m}^{2}$. Orange: $A=20\times 20\,\mathrm{\mu m}^{2}$. Green: $A=30\times 30\,\mathrm{\mu m}^{2}$. Circles mark the pull-in instability locations. The etch depth is $d=500\,\mathrm{nm}$. The effective spring constant was calculated analytically from Eq. (26) assuming a $30\,\mathrm{\mu}m$ silicon nitride square membrane vibrating in the fundamental mode with $\sigma=1\,\mathrm{GPa}$ and $\rho=3200\,\mathrm{kg\cdot m^{-3}}$.
@@ VI Evanescent fields
So far in this paper we have separately introduced membrane waveguides and resonators, along with the continuum and lumped-element models. We now shift focus towards how these circuit elements can be coupled. Coupling is an essential ingredient in phononic circuits; we will show that one of the outstanding aspects of suspended membranes is that the coupling can be engineered very flexibly and precisely. We will build towards this conclusion in the rest of the paper and finish by outlining several examples. Firstly we introduce evanescent acoustic fields, which underpin the mechanism by which acoustic phonons are coupled.
@@
Evanescent waves are also very useful for membrane phononic circuits, where they allow single mode waveguides and highly efficient coupling of heterogeneous membrane geometries [3, 22, 58]. A simple use case for these functions is the coupling of a resonator with input and output single mode waveguides, allowing mechanical logic to be performed without electromechanical actuation [20]. In this tutorial we will go further and detail several more phononic devices that can be achieved using evanescent couplers (see Section XI).
To see how evanescent waves arise in membrane phononics, consider a waveguide of width $W$ defined by having a density $\rho_{1}$ greater than the density $\rho_{2}$ of the surrounding material, as illustrated in Fig. 20. This is not the same mechanism of operation as membrane waveguides defined by the large impedance mismatch between the membrane and substrate (described earlier), but it has the same principle and makes for a simpler example.
@@
$\psi(x)=\begin{cases}\cos(k_{x}x)&|x|<\frac{W}{2}\\ e^{-\kappa(|x|-\frac{W}{2})}&\text{elsewhere}.\end{cases}$ (71)
@@
$\frac{\rho_{2}\Omega^{2}}{\sigma}$ $=k_{y}^{2}-\kappa^{2}.$ (74)
Combining Eqs. (72), (73), and (74) eliminates $k_{y}$ and $\kappa$, yielding a transcendental equation for $k_{x}$:
@@
Figure 21: Equivalence between confinement of optical waves with dielectric index contrast and confinement of acoustic waves with density contrast. (a) In a dielectric slab or optical fiber, the guided mode is confined into the core, which has high refractive index than cladding material. The mode evanescently extends into the cladding, but no energy is transferred outside, such that is can analogously be formulated in a ray optics description in terms of total internal reflection (dashed line). (b) A mathematically equivalent example is the phononic waveguide studied in this section, where density differences create a guided mode.
Equation (71) is mathematically equivalent to the case of the transverse electric mode of light in a dielectric waveguide, if we make the substitution $\rho\mapsto\varepsilon$ [63]; that is, if we replace density differences with electric permittivity differences. This equivalence is illustrated in Fig. 21. Because refractive index scales with $\varepsilon^{2}$, the density contrast of $\rho_{1}/\rho_{2}=4$ illustrated in Fig. 20 is equivalent to a refractive index contrast of $n_{2}/n_{1}=2$, such as is found for silicon nitride photonic waveguides surrounded by air [126, 127].